OpenStudy (anonymous):

Use the distributive property to simplify the expression below: 4x-10/2

OpenStudy (anonymous):

\[\frac{4x - 10}{2} \implies \frac{2(2x-5)}{2} \implies \frac{\cancel{2}(2x - 5)}{\cancel{2}} \implies (2x - 5)\]

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

yeah its correct

OpenStudy (anonymous):

okay thx!

OpenStudy (anonymous):

Yeah it is right..